The finishing process for automotive body panels has two main steps: painting and clear coat. The times to complete each step are Normally distributed random variables. The painting step takes an average of 25 minutes, with a standard deviation of 3 minutes. The clear coat step takes an average of 17 minutes, with a standard deviation of 2 minutes.

(a) What is the probability that a body panel with be finished within 35 minutes?

(b) What is the probability that it will take longer than 40 minutes?
Answer #1

expected amount for total panel =25+17=42

and std deviation =sqrt(32+22)= 3.6056

a)

probability that a body panel with be finished within 35 minutes =P(X<35)=P(Z<(35-42)/3.6056)=P(Z< -1.94) =0.0262

b)

probability that it will take longer than 40 minutes =P(X>40)=P(Z>(40-42)/3.6056)=P(Z> -0.55)=0.7088

( please try 0.7106 if this comes wrong due to rounding error)